OpenStudy (anonymous):

factor (x^2-1)x^4n-(x^2y^8n-y^8n) completely.

OpenStudy (smileyxl3):

\[-nx^2y^8+ny^8+nx^6-nx^4\]

OpenStudy (mathstudent55):

Is the n of 4n part of the exponent or not?

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

Is this the problem where 8n is an exponent? \( (x^2-1)x^{4n} - (x^2y^{8n}-y^{8n}) \)

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

\( (x^2-1)x^{4n} - (x^2y^{8n}-y^{8n})\) Factor out y^(8n) from the second part. \( =(x^2 - 1)x^{4n} - y^{8n} (x^2 - 1) \) Now you have a commonn factor of x^2 - 1

OpenStudy (mathstudent55):

\(=(x^2 - 1)(x^{4n} - y^{8n}) \) Now use difference of squares on each factor.

OpenStudy (anonymous):

so is it (x+1)(x-1)1^4n(x-y^2n) ?

OpenStudy (mathstudent55):

\( (x + 1)(x - 1)(x^{2n} + y ^{4n}) (x^{2n} - y^{4n} ) \) The last term above can be factored again using difference of squares.

OpenStudy (anonymous):

okay thank you

OpenStudy (mathstudent55):

\( = (x + 1)(x - 1)(x^{2n} + y ^{4n}) (x^n + y^{2n} ) (x^n - y^{2n} )\)

OpenStudy (mathstudent55):

wlcm